OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2). I got the answer y=3x-5 If I am wrong, well you walk me through it?

jimthompson5910 (jim_thompson5910):

that's correct, congrats

OpenStudy (anonymous):

Thanks(:

jimthompson5910 (jim_thompson5910):

to check, notice how when x = 1, y is.. y = 3x-5 y = 3(1)-5 y=3-5 y=-2 So when x =1, y = -2 giving the point (1, -2)

jimthompson5910 (jim_thompson5910):

np